Two charges pellets, each with a charge of 5C, are located a distance 15m apart. What is the electric force between the two pellets?
1
Expert's answer
2019-11-21T09:59:09-0500

Coulomb's law is a physical law that describes the force of interaction between fixed point electric charges depending on the distance between them.

In its scalar form, the law is:

"F=k\frac{|q_1||q_2|}{r^2}, \space where \space k=8.9876\cdot10^9\space \frac{N\space m^2 }{C^2}"

So, two pellets have the same sign of charge, therefore the repulsive force

"F=8.9876\cdot10^9\space \frac{N\space m^2 }{C^2} \space \frac{5^2\space C^2}{15^2\space m^2}\approxeq 10^9\space N"

Answer: the magnitude of the electric force between two pellets is "10^9\space N"